OpenStudy (anonymous):

Find the equation of a line through (3,6) and parallel to y=4x-1

OpenStudy (anonymous):

parallel means that you keep the same slope but need to change the y intercept.

OpenStudy (anonymous):

The equation would be y=4x-6 because if it is parallel to your equation it has the same slope so all you have to use is y-y1=m(x-x1) so it is y-6=4(x-3) all you have to do then is solve for y which =4x-6